### v1

# DesignX: Human-Competitive Algorithm Designer for Black-Box Optimization

Abstract
Designing effective black-box optimizers is hampered by limited problem-specific knowledge and manual control that spans months for almost every detail. In this paper, we present DesignX, the first automated algorithm design framework that generates an effective optimizer specific to a given black-box optimization problem within seconds. Rooted in the first principles, we identify two key sub-tasks: 1) algorithm structure generation and 2) hyperparameter control. To enable systematic construction, a comprehensive modular algorithmic space is first built, embracing hundreds of algorithm components collected from decades of research. We then introduce a dual-agent reinforcement learning system that collaborates on structural and parametric design through a novel cooperative training objective, enabling large-scale meta-training across 10k diverse instances. Remarkably, through days of autonomous learning, the DesignX-generated optimizers continuously surpass human-crafted optimizers by orders of magnitude, either on synthetic testbed or on realistic optimization scenarios such as Protein-docking, AutoML and UAV path planning. Further in-depth analysis reveals DesignX’s capability to discover non-trivial algorithm patterns beyond expert intuition, which, conversely, provides valuable design insights for the optimization community. We provide DesignX’s inference code at https://github.com/MetaEvo/DesignX.
1 Introduction
Black-box optimization (BBO) lies at the core of scientific and industrial advances, such as electronic design automation [1], molecular design [2] and AutoML [3]. Yet, BBO is challenging due to unavailable objectives and derivatives, and complex, diverse properties that demand extensive expert knowledge. Evolutionary Computation (EC) is widely recognized as a robust derivative-free paradigm for BBO [4]. Since the 1990s, numerous EC variants such as genetic algorithms[5], differential evolution [6], particle swarm optimization [7], and evolution strategies [8] have emerged. Despite shared core paradigm, they rely on expert-designed adaptive operators [9] and hyperparameter control [10] to achieve the best performance on a particular BBO class or instance.
However, manually redesigning optimizers for each new BBO problem is neither scalable nor practical. Recently, an emerging research avenue termed as Meta-Black-Box-Optimization (MetaBBO) [11] has emerged, which automates algorithm design (AAD) through a bi-level paradigm: a meta-level learns a policy to guide low-level BBO optimizer. By meta-training [12] over a distribution of problems, MetaBBO can generate customized algorithms for both seen and unseen instances.
Despite the success, existing MetaBBO approaches merely focus on learning specific sub-tasks of AAD for EC. Specifically, optimizer design involves two sequential sub-tasks (see Figure 1, top left): (1) determining the algorithm workflow, and (2) control its internal hyperparameters. Existing work addresses the former via algorithm selectors [13, 14, 15] or workflow generators [16, 17, 18, 19], and the latter through reinforcement learning (RL) [20] for online control [21, 22, 23, 24]. While learning a single sub-task eases training, it often results in sub-optimal designs and limits potential performance gains.
Figure 1: Left: Compared to manual design process, DesignX replaces human experts by two learnable agents. Right: Four dashed lines denote average performances of well-known human-crafted optimizers in decades. During pre-training, DesignX surprisingly discovers powerful optimizers superior to the ones crafted by human experts.
In this paper, we advance MetaBBO research by proposing the first unified framework that jointly learns both sub-tasks of algorithm design—workflow generation and hyperparameter control, so as to enable the discovery of human-competitive optimizers in an end-to-end fashion.
This is achieved through several key innovations. Firstly, we extend and enrich the Modular-BBO modularization system in [24], resulting in a more comprehensive system: Modular-EC. Specifically, since Modular-BBO is primarily constructed for DE optimzier, Modular-EC integrates more diverse sub-modules in ES, GA and PSO into its sub-module library. Modular-EC now supports representing different optimizer types, enhancing the capacity of Modular-BBO. Building on the upgraded Modular-EC, we develop a dual-agent reinforcement learning system (see Figure 1, bottom left), where both agents are Transformer-based [25]: 1) Agent-1 autoregressively samples valid optimizer workflows conditioned on the problem instance; 2) Agent-2 dynamically adjusts hyperparameters during optimization by incorporating real-time feedback. A novel cooperative reward scheme encourages both agents to make mutually conditioned decisions, jointly optimizing for maximum performance. We train this dual-agent system on a large-scale problem set of 10k synthetic instances, and observe it consistently discovering optimizers that outperform expert-crafted baselines (see Figure 1, right). Remarkably, through days of autonomous learning, the DesignX-generated optimizers continuously surpass human-crafted optimizers by orders of magnitude, either on synthetic testbed or on realistic optimization scenarios such as Protein-docking, AutoML and UAV. Furthermore, the testing results clearly demonstrate the novelty and superiority of DesignX against up-to-date MetaBBO baselines. To summarize, the contributions of this paper are in three folds:
  • The first MetaBBO framework that achieves fully end-to-end AAD for BBO problems, paving the way of developing foundation model in this domain.
  • We obtain a well-performing model (DesignX) through large-scale training, capable of designing powerful optimizers for diverse, unseen, realistic problems
  • Further in-depth analysis reveals the importance of the proposed novel designs, providing first-hand insights on non-trivial algorithm patterns beyond expert intuition.
2 Related Works
We review the development of Automated Algorithm Design (AAD) over the past decades. Early efforts by Schmidhuber et al. [26] applied Genetic Programming (GP) to recursively improve another GP in a self-referential manner. Later, GP was applied to design full algorithm templates [27], but difficulties in genotype design and expensive evaluations limited its scalability for BBO problems. Recent MetaBBO approaches integrate machine learning techniques such as reinforcement learning (RL) and large language models (LLMs) to develop more flexible and generalizable optimizers [28, 11]. These RL-based methods like DEDQN [14] and DEDDQN [13] focused on operator selection and hyperparameter control within fixed algorithm structures. More recent methods leverage Transformer architectures for enhanced control [23, 29], including ConfigX [24] and Q-Mamba [30], which implement online and offline RL, respectively. Other works explored Transformer-based generation of algorithm components. SYMBOL [31] learned to compose new operators as symbolic sequences. ALDes [17] tokenized common algorithmic modules and turned workflow design into sequence generation. GLHF [32] simulated DE operators with trainable modules optimized through gradient descent. Though these models were relatively small, they achieved strong performance. With LLM-scale models, capabilities expand further. LLMs can search reward functions [33], optimize neural architectures [34], act as optimizers based on previous search trajectories [35], or generate algorithm code from problem descriptions [36, 19, 18]. However, existing work focuses on only one sub-task of AAD: either generating workflows or controlling parameters. No prior method jointly addresses both, which motivates our proposed DesignX to enable end-to-end algorithm design.
3 Methodology
3.1 Modular-EC
Existing EC optimizers commonly comprise a series of algorithm modules. A massive array of novel algorithm modules have been proposed in literature for specific optimization scenarios [9, 37, 38]. It is a quite natural idea to “stand on the shoulder of giants” for designing new optimizers, that is to say, construct a modular algorithmic space and search for well-performing optimizer workflow in it [39, 40]. Following such idea, ConfigX [24] proposes a comprehensive modularization system: Modular-BBO for learning universal hyper-parameter control policy in DE. It groups commonly used sub-module variants in existing DE optimizers into 9 module types: 6 of which are Uncontrollable without hyper-parameters: Initialization [41], Boundary_Control [42], Selection [43], Niching [44], Restart_Strategy [45], Population_Reduction [46], and the rest 3 of which are Controllable with hyper-parameters: Mutation [47], Crossover [48], and Information_Sharing [49].
In Modular-EC, we have added a novel module type Other_Update [50, 51] into Modular-BBO’s module library, which belongs to Controllable namespace. We integrate popular reproduction operators of diverse ES, GA and PSO optimizers into Other_Update and also update the other 9 module types by adding corresponding sub-modules in ES, GA and PSO. To summarize, Modular-EC supports 10 module types with 116 module variants in total. This results in millions of possible algorithm workflows, significantly enhancing the expressiveness of Modular-BBO.
For a concrete module variant, Modular-EC assigns it an unique $16$-bit binary code id for identify. A topology_rule list is built within each module variant to indicate which module types are allowed to be placed right after this module variant, ensuring legal generation of optimizer workflow in auto-regressive fashion. We list some examples here: 1) Any EC optimizer must start with Initialization; 2) Boundary_Control is not allowed placed between two subsequent reproduction modules (e.g., Mutation and Crossover); 3) Restart_Strategy is only allowed to be placed at the end of a EC optimizer. We provide more details of the hierarchical architecture, module variants information of Modular-EC in Appendix A.
3.2 Dual-agent Algorithm Design System
We propose a dual-agent algorithm design system for DesignX to operate on Modular-EC. As shown in Figure 2, the system consists of two Transformer-based RL agents: Agent-1 ($\pi_{\phi}$) and Agent-2 ($\pi_{\theta}$), each addressing a core sub-task in automated algorithm design. 1) Algorithm workflow generation: Agent-1 constructs a customized optimizer workflow based on the given problem. 2) Hyperparameter control: Agent-2 dynamically adjusts the hyperparameters during the optimization process to enhance performance. By jointly addressing both sub-tasks, DesignX offers a more complete and effective solution than methods focusing on only one aspect (see Section 2).
Before we get into further technical details, we first explain the Program Structure Tree (PST) [52] of an algorithm workflow and pre-order traversal of PST. We illustrate a simple example in the left of Figure 2, where a two-population niching-based EC optimizer is represented by PST and corresponding pre-order traversal respectively. The pre-order traversal representation of an optimizer workflow is primarily used in Agent-1 and Agent-2 to align with information processing logic of Transformer architecture, where each module in the traversal is regarded as a token.
3.2.1 Agent-1: Determine the Workflow
Agent-1’s workflow is shown in the top right of Figure 2. Given the feature vector $\mathcal{F}_{p}$ of an optimization problem $p$, Agent-1 auto-regressively samples module variants from Modular-EC to construct a complete optimizer workflow $\mathcal{A}p=\pi\phi(\mathcal{F}p)$. The architecture of $\pi\phi$ consists of four components: 1) a problem feature embedder $\mathcal{W}_{token}\in\mathbb{R}^{13\times h}$, where 13 is $\mathcal{F}_{p}$’s dimension and $h$ denotes the token embedding dimension; 2) a Tokenizer $\mathcal{W}_{token}\in\mathbb{R}^{16\times h}$, where 16 denotes the 16 bits module id; 3) $L$ sequential GPT-2 [53] blocks with $k$ heads and hidden dimension $h$. We use $MSA_{1}$ to denote these attention blocks; 4) a masked Softmax module $\mathcal{W}_{sample}\in\mathbb{R}^{h\times 117}$, where 117 is the numbers of tokens (116 modules in Modular-EC and an additional end token to terminate generation).
Problem Feature Embedding. The raw feature $\mathcal{F}_{p}$ for a given optimization problem $p$ is a 13-dimensional vector, which is further divided into two parts: 1) 4 basic properties: the dimension, allowed maximum function evaluations, upperbound and lowerbound of searching range; 2) 9 statistical properties: we use a well-known optimization problem statistical analysis framework, Exploratory Landscape Analysis (ELA) [54], which provides many statistical low-level features for profiling high-level optimization properties such as multi-modality, separability, global structure, etc. Specifically, we select 9 ELA features with both significant independence and efficient computation according to the sensitivity analysis of ELA features in [55, 56]. We provide a detailed elaboration on these ELA features in Appendix B.1. Once $\mathcal{F}_{p}$ is obtained, we use $\mathcal{W}_{token}$ to map it to a $h$-dimensional token, which we denote as start for subsequent optimizer workflow generation.
Figure 2: Left: The dual-agent system in DesignX processes an optimizer workflow by the pre-order traversal of its program structure tree. Top Right: Agent-1 generates legal optimizer workflow in an auto-regressive fashion. Bottom Right: Agent-2 controls hyperparameters of the generated optimizer workflow by conditioning on the optimization progress information.
Auto-regressive Generation. Starting from the start token for problem $p$, Agent-1 auto-regressively generates the pre-order traversal of an optimizer workflow $\mathcal{A}_{p}$. Suppose Agent-1 has generated $m$ modules $\{\mathcal{A}_{p}^{1},\mathcal{A}_{p}^{2}...,\mathcal{A}_{p}^{m}\}$, then the sampling distribution of $(m+1)$-th module $\mathcal{A}_{p}^{m+1}$ is:
$P(\mathcal{A}_{p}^{m+1}|start,\mathcal{A}_{p}^{1},...,\mathcal{A}_{p}^{m})\sim Softmax(mask(\mathcal{A}_{p}^{m})\odot(\mathcal{W}_{sample}^{\mathrm{T}}\cdot H^{(m)})),$ (1)
$H=MSA_{1}(Pos+\{start,\mathcal{W}_{token}^{\mathrm{T}}\cdot\mathcal{A}_{p}^{1}.get\_id(),...,\mathcal{W}_{token}^{\mathrm{T}}\cdot\mathcal{A}_{p}^{m}.get\_id()\})$
where we first get each sampled module’s id and use the tokenizer to map them to tokens with $h$-dimension. Then all tokens including start are added with Cosine Position Encoding $Pos$. After going through the GPT-2 blocks $MSA_{1}$, we use $\mathcal{W}_{sample}$ to map the output embedding $H^{(m)}$ for $m$-th module as the prediction head. Recall that we have to ensure the generated workflow is legal. To achieve this, we propose a masked Softmax sampling procedure. A boolean mask vector $mask(\mathcal{A}_{p}^{m})\in\mathbb{R}^{117}$ is obtained by checking $\mathcal{A}_{p}^{m}$’s topology rule $\mathcal{A}_{p}^{m}.get\_rule()$. Hadamard product between the mask and prediction head squeezes the sampling probability of illegal modules to 0. We note that the dimension of prediction head and the mask is 117, which corresponds to the 116 modules in Modular-EC and the end token. Without the end token, Agent-1 has risks of generating infinite trajectory. Refer to Appendix A, Table 2 to check which modules could be placed right before $end$. In the rest of this paper, we use $\pi_{\phi}(\mathcal{A}_{p})$ to denote the sampling probability of a concrete workflow $\mathcal{A}_{p}$, which is the successive multiplication of all generation steps:
$\pi_{\phi}(\mathcal{A}_{p})=P(\mathcal{A}_{p}^{1}|start)P(\mathcal{A}_{p}^{2}|start,\mathcal{A}_{p}^{1})...P(end|start,\mathcal{A}_{p}^{1},...,\mathcal{A}_{p}^{M})$ (2)
3.2.2 Agent-2: Control the Hyper-parameters
Agent-2’s workflow is shown in the bottom right of Figure 2. Once $\mathcal{A}_{p}$ is generated by Agent-1, it is used to optimize $p$. During the optimization process, given some observed optimization progress information $\mathcal{O}_{t}$ at $t$-th optimization step, Agent-2 dynamically adjusts hyper-parameter values $\mathcal{C}_{t}=\pi_{\theta}(\mathcal{O}_{t})$ for all Controllable modules in $\mathcal{A}_{p}$. The motivation behind Agent-2 is that: a common observation in EC domain reveals that hyperparameter values in an optimizer more or less impact the exploration/exploitation tradeoff [9]. An effective parameter control policy could further enhance the optimization performance of the optimizer generated by Agent-1.
To suggest per optimization step hyperparameter values for Controllable modules in $\mathcal{A}_{p}$, an informative optimization progress feature vector $\mathcal{O}_{t}$ is first computed following the common idea of up-to-date MetaBBO approaches [23, 31, 24]. $\mathcal{O}_{t}$ is a 9-dimensional vector of which each dimension is a statistical feature indicating the local/global distribution in solution/objective space, convergence progress and optimization budget usage information. We provide detailed description of these features in Appendix B.2. Agent-2 then embeds $\mathcal{O}_{t}$ into each module in $\mathcal{A}_{p}$ to get all module’s embeddings:
$Emb(\mathcal{A}_{p}^{m})=Pos+\mathcal{W}_{emb}^{\mathrm{T}}\cdot[\mathcal{A}_{p}^{m}.get\_id(),\mathcal{O}_{t}]\quad m=1,2...,M$ (3)
where $\mathcal{W}_{emb}^{\mathrm{T}}\in\mathbb{R}^{25\times h}$ maps the concat of module id and $\mathcal{O}_{t}$ to $h$-dimensional embeddings. The final embedding for each module is obtained by adding the $h$-dimensional embeddings with Cosine Positional Embedding codes, which inject relative order information among the modules to let Agent-2 aware of the optimizer workflow structure. The suggested hyperparameter values at optimization step $t$ is decoded by first feeding the embeddings of all modules into $L$ sequential GPT-2 [53] blocks with $k$ heads and hidden dimension $h$ (denoted as $MSA_{2}$). Then the output decision embeddings $H_{dec}$ are further decoded into normal distribution parameters:
$\mu=\mathcal{W}_{\mu}^{\mathrm{T}}\cdot H_{dec},\quad\Sigma=\mathcal{W}_{\Sigma}^{\mathrm{T}}\cdot H_{dec},\quad H_{dec}=MSA_{2}(Emb(\mathcal{A}_{p}^{1}),...,Emb(\mathcal{A}_{p}^{M}))$ (4)
where $\mathcal{W}_{\mu}^{\mathrm{T}}\in\mathbb{R}^{h\times N_{max}}$ and $\mathcal{W}_{\mu}^{\mathrm{T}}\in\mathbb{R}^{h\times N_{max}}$ are network parameters of the hyperparameter value head. They map $H_{dec}$ to the mean parameters $\mu\in\mathbb{R}^{M\times N_{max}}$ and covariance parameters $\Sigma\in\mathbb{R}^{M\times N_{max}}$, where $\mu^{(m)}\in\mathbb{R}^{N_{max}}$ and $\Sigma^{(m)}\in\mathbb{R}^{N_{max}}$ denotes distribution parameters for $m$-th module in $\mathcal{A}_{p}$. At last, the hyperparameter values $C_{t}$ are sampled from the predicted normal distributions for all $M$ modules:
$C_{t}=\{C_{t}^{1},...,C_{t}^{M}\}\sim\{\mathcal{N}(\mu^{(1)},\Sigma^{(1)}),...,\mathcal{N}(\mu^{(M)},\Sigma^{(M)})\}$ (5)
We have to note that since different modules in Modular-EC might hold different number of hyperparameter values, we predefine a maximum configuration size $N_{max}$ to cover them. If the number of hyper-parameters in a module is less then $N_{max}$, we use the first few sampled values and ignore the rest. Suppose the optimization horizon for problem $p$ is $T$ steps, Agent-2 will be asked $T$ times for deciding the per-step hyper-parameter values. In the rest of this paper, we use $\pi_{\theta}(C_{t}|\mathcal{A}_{p})$ to denote the associate probability of the hyperparameters for $\mathcal{A}_{p}$ at optimization step $t$ [^1] We only consider sampling for modules with at least one hyperparameter.:
$\pi_{\theta}(C_{t}|\mathcal{A}_{p})=\prod_{m=1}^{M}\mathcal{N}(\mu^{(m)},\Sigma^{(m)})$ (6)
3.3 Cooperative Large Scale Training
We propose a large scale meta-reinforcement-learning paradigm to ensure the pre-trained DesignX model could benefit from the harmonious cooperation between Agent-1 & 2, and is capable of being generalized towards unseen problems.
Large Scale Synthetic Problem Set. We construct a large scale synthetic problem set containing 12800 diverse problem instances for the ease of training generalizable DesignX model. 32 representative basic problems are first collected from popular BBO benchmarks [57, 58], including Rastrigin, Schwefel, Rosenbrock, etc. We follow the steps below to generate 12800 diverse problem instances: 1) We first define three problem construction modes, “single”, “composition” and “hybrid”. “single” mode randomly selects one basic problem. “composition” mode randomly aggregates 2-5 basic problems by weighted summation of their objective functions. “hybrid” mode divides decision variables into some subcomponents and then randomly selects a group of basic functions, which are used for different subcomponents. 2) By randomly selecting the construction modes and determining the searching range, dimension (5-50d), maximum allowed optimization budget (10000-50000 maxFEs) and rotation/shift in solution space, we construct 12800 problem instances with diverse optimization properties, which aligns with the intricate problem distribution in real world. We further randomly split them into a training problem set $\mathcal{D}_{train}$ (9600 instances) and a testing set $\mathcal{D}_{test}$ (3200 instances). A more detailed elaboration is provided in Appendix C.
Cooperative Training Objective. We formulate the automated algorithm design task of DesignX as a dual-agent Markov Decision Process (MDP). For each problem instance $p\in\mathcal{D}_{train}$, Agent-1 first generates a legal optimizer workflow $\mathcal{A}_{p}$ with probability $\pi_{\phi}(\mathcal{A}_{p})$ in Eq. (2). $\mathcal{A}_{p}$ is then used to optimize $p$ until its allowed optimization budget is used up. For each optimization step $t$ along this optimization process ($T$ steps in total), Agent-2 continuously dictates hyperparameters $C_{t}$ with probability $\pi_{\theta}(C_{t}|\mathcal{A}_{p})$ in Eq. (6). We record the reward obtained at $t$-th step as $r_{t}=\frac{f_{p}^{t-1,*}-f_{p}^{t,*}}{f_{p}^{0,*}-f_{p}^{*}}$, where $f_{p}^{t,*}$ denotes the optimal objective value found until $t$-th step (w.l.o.g., $p$ is assumed as a minimization problem), $f_{p}^{*}$ denotes the optimal objective value of $p$. Then the training objective of DesignX’s MDP can be formulated as:
$\mathcal{J}(\phi,\theta)=\mathbb{E}_{p\sim\mathcal{D}_{train}}[\sum_{t=1}^{T}r_{t}]=\frac{1}{|\mathcal{D}_{train}|}\sum_{i=1}^{|\mathcal{D}_{train}|}\sum_{t=1}^{T}r_{t}$ (7)
which is the expected optimization performance if we use DesignX’s Agent-1 & 2 to design optimizers for solving problem instances in $\mathcal{D}_{train}$. For Agent-1, there is no intermediate reward (delayed-reinforcement task), hence we train it by episodic reinforcement learning method REINFORCE [59]. For Agent-2, the per-step reward $r_{t}$ can be used hence we train it by the popular PPO method [60]. We provide the pseudo code of the training procedure in Appendix D, Alg. 1.
4 Experimental Analysis
[…]
Experiments Setup. The baselines in experiments include: 1) a DesignX model trained after 6 days; 2) up-to-date MetaBBO approaches GLHF [32], DEDQN [14] and GLEET [23] that excel at workflow learning or hyper-parameter control; 3) representative human-crafted optimizers: a) those before 2000, GA [5], PSO [7] and DE [6]. b) those in 2000-2010, CMAES [61], FIPSO [62], SaDE [63], CLPSO [64] and JADE [65]. c) those in 2010-2020, CoDE [66], IPSO [67], SHADE [68], LM-CMA-ES [69] and GLPSO [70]. d) those after 2020, MadDE [71], jDE21 [72], MMES [73] and NL-SHADE-LBC [74]. For evaluation fairness, we train DesignX and other MetaBBO baselines on the same $\mathcal{D}_{train}$ (see Section 3.3). We leave detailed training settings and other hyper-parameter settings of all baselines at Appendix E.1 & E.2. To simplify presentation, we use following tags: “MetaBBO”, “’before 00’, “00s”, “10s” and “after 20” to tag these baseline.
4.1 Performance Comparison (RQ1)
Table 1: The in-distribution generalization performance in terms of absolute optimization performance results on $\mathcal{D}_{test}$. The best is labeled in green and the second best is labeled in red.
before 00 00s 10s after 20 MetaBBO DesignX
F1
MAH, 50D, 30000 FEs
6.60E+00
$\pm$3.74E+00
1.64E+00
$\pm$1.64E+00
1.27E+00
$\pm$4.41E-01
5.32E+00
$\pm$3.70E+00
2.80E+00
$\pm$0.00E+00
2.89E-01
$\pm$3.93E-01
F79
UAH, 5D, 50000 FEs
2.98E+00
$\pm$9.95E-01
3.70E+00
$\pm$1.71E+00
5.38E+00
$\pm$4.05E-01
1.81E+00
$\pm$1.83E-01
9.95E-01
$\pm$0.00E+00
5.68E-02
$\pm$1.17E+00
F125
UAH, 10D, 40000 FEs
1.39E-03
$\pm$1.38E-03
3.50E-06
$\pm$3.50E-06
1.48E-04
$\pm$1.33E-04
1.69E-05
$\pm$7.99E-06
1.08E-04
$\pm$0.00E+00
4.81E-07
$\pm$2.66E-07
F154
UAH, 50D, 10000 FEs
1.35E+03
$\pm$2.26E+02
1.44E+03
$\pm$3.45E+02
1.38E+03
$\pm$2.40E+02
1.46E+03
$\pm$6.17E+02
5.47E+02
$\pm$0.00E+00
6.99E+02
$\pm$7.45E+01
F211
MAH, 5D, 40000 FEs
6.55E-01
$\pm$2.92E-01
8.04E-01
$\pm$6.99E-01
2.64E-01
$\pm$9.96E-02
1.28E-01
$\pm$2.43E-02
1.59E-01
$\pm$0.00E+00
7.28E-02
$\pm$6.56E-02
F240
MWL, 20D, 20000 FEs
6.39E+00
$\pm$4.25E+00
8.72E+00
$\pm$1.71E+00
8.24E+00
$\pm$2.19E+00
3.97E+00
$\pm$3.31E+00
2.05E+00
$\pm$0.00E+00
1.27E-01
$\pm$2.99E+00
F326
UAL, 10D, 40000 FEs
1.10E+00
$\pm$1.22E-01
1.15E+00
$\pm$4.53E-01
2.47E+00
$\pm$4.64E-01
7.66E-01
$\pm$5.10E-02
1.22E+00
$\pm$0.00E+00
5.84E-01
$\pm$1.66E+00
F411
UAL, 10D, 50000 FEs
2.50E-01
$\pm$9.51E-02
4.07E-01
$\pm$9.27E-02
2.68E-01
$\pm$7.41E-02
1.28E-01
$\pm$1.11E-02
1.87E-01
$\pm$0.00E+00
7.91E-02
$\pm$4.83E-02
F545
UWL, 5D, 40000 FEs
2.98E+00
$\pm$9.94E-01
1.49E+00
$\pm$4.97E-01
3.36E+00
$\pm$6.23E-01
8.41E-01
$\pm$1.54E-01
1.99E+00
$\pm$0.00E+00
2.61E-08
$\pm$6.48E-01
F1045
MWH, 10D, 40000 FEs
7.53E+02
$\pm$1.68E+00
4.20E+02
$\pm$1.47E+02
2.29E+02
$\pm$6.06E+01
1.71E+02
$\pm$1.41E+01
9.21E+02
$\pm$0.00E+00
1.67E+02
$\pm$1.06E+02
F1139
MAH, 10D, 50000 FEs
1.16E+01
$\pm$1.00E+01
3.95E+00
$\pm$1.67E+00
8.90E+00
$\pm$1.54E+00
2.67E+00
$\pm$1.15E+00
1.19E+01
$\pm$0.00E+00
1.39E-04
$\pm$1.47E+00
F1200
MAL, 50D, 40000 FEs
7.47E+00
$\pm$6.29E+00
1.14E+00
$\pm$7.00E-03
1.15E+00
$\pm$2.56E-02
1.33E+01
$\pm$1.22E+01
1.27E+00
$\pm$0.00E+00
1.09E+00
$\pm$2.12E-02
F1556
MAH, 10D, 40000 FEs
3.99E+02
$\pm$3.75E+02
2.03E+02
$\pm$1.76E+02
2.73E+01
$\pm$7.63E+00
1.48E+01
$\pm$7.13E-01
9.45E+00
$\pm$0.00E+00
1.01E+01
$\pm$1.13E+02
F1653
MAH, 20D, 10000 FEs
2.55E+01
$\pm$1.53E+00
2.53E+01
$\pm$7.11E-01
2.72E+01
$\pm$7.28E-01
1.85E+01
$\pm$2.56E+00
1.69E+01
$\pm$0.00E+00
1.54E+01
$\pm$3.47E+00
F1687
MAL, 50D, 40000 FEs
8.98E+00
$\pm$6.60E+00
2.07E+01
$\pm$1.07E+01
4.49E-01
$\pm$2.38E-01
2.94E+01
$\pm$2.81E+01
1.94E+00
$\pm$0.00E+00
2.24E-02
$\pm$9.09E+00
F2068
MWH, 20D, 20000 FEs
3.79E+01
$\pm$6.53E+00
2.32E+00
$\pm$1.13E-01
1.46E+01
$\pm$1.40E+01
1.65E+01
$\pm$1.41E+01
3.72E+01
$\pm$0.00E+00
5.16E-01
$\pm$1.06E+01
F2390
MAL, 10D, 30000 FEs
3.93E+00
$\pm$2.15E+00
2.78E+00
$\pm$0.00E+00
6.34E+00
$\pm$9.03E-01
1.54E+00
$\pm$1.10E+00
2.04E+01
$\pm$0.00E+00
1.85E-03
$\pm$2.45E+00
F2473
MAL, 10D, 20000 FEs
1.10E+00
$\pm$9.06E-02
3.98E-01
$\pm$6.88E-02
8.72E-01
$\pm$1.80E-02
6.69E-01
$\pm$2.53E-01
1.42E-01
$\pm$0.00E+00
1.63E-01
$\pm$1.59E-01
F2895
MWL, 10D, 50000 FEs
1.90E+01
$\pm$3.88E+00
4.34E+00
$\pm$1.66E+00
1.18E+01
$\pm$5.35E+00
4.23E+00
$\pm$5.26E-01
4.98E+00
$\pm$0.00E+00
1.99E+00
$\pm$3.34E+00
F2986
MAL, 50D, 10000 FEs
4.37E+02
$\pm$1.71E+02
4.93E+02
$\pm$3.74E+02
1.60E+02
$\pm$6.45E+01
2.51E+03
$\pm$2.42E+03
1.01E+02
$\pm$0.00E+00
8.90E+01
$\pm$2.65E+01
Normalized Averaged
Objective
2.94E-01
$\pm$1.01E+00
1.96E-01
$\pm$1.62E+00
1.54E-01
$\pm$2.61E-01
1.46E-01
$\pm$2.35E-01
1.32E-01
$\pm$7.36E-01
8.26E-02
$\pm$1.75E-01
In-distribution Generalization. All baselines are tested on our proposed $\mathcal{D}_{test}$ (see Section 3.3), with 51 independent runs for each problem instance. Due to the space limitation, we present the absolute optimization performance of all baselines on 20 of the 3200 tested instances in Table 1. These 20 instances are randomly selected to showcase their diversity in: a) optimization properties, “U/M” for unimodal/multi-modal, “A/W” for adequate or weak global structures and “L/H” for low or high conditioning; b) problem dimensions, 5D-50D; c) allowed optimization budget in terms of the number function evaluations (FEs). We additionally average the baselines in each tag (“before 00”, “00s” and etc.) for the ease of presentation. Refer to this link for complete results of each baseline across all 3200 problem instances.
The results in Table 1 reveal that: 1) The human-crafted BBO optimizers achieve progressive advancement through the expert-level designs proposed over the past decades. However, they are sill restricted by no-free-lunch theorem. 2) By incorporating learning paradigm into BBO optimizers, MetaBBO approaches are capable of boosting the low-level optimizers on some problem instances. 3) The optimization performance of DesignX surpasses both MetaBBO and hand-crafted BBO baselines, ranking the first place on almost all tested instances with diverse properties. Through learning the bi-agent system across a large scale problem distribution ($\mathcal{D}_{train}$), DesignX intelligently designs powerful and customized optimizers for different problems. To the best of our knowledge, this is the first time a RL system successfully learns how to automatically design BBO optimizers.
Figure 3: The generalization performance of baselines on realistic scenarios.
Out-of-distribution Generalization. For learning-assisted optimization techniques, the problem shifts in realistic scenarios might challenge their generalization ability in practice. To this end, we test DesignX and MetaBBO baselines trained on synthetic $\mathcal{D}_{train}$ on three diverse realistic BBO testsuites: a) Protein-Docking [75], a collection of 280 protein-docking instances, featured by intricate landscapes; b) HPO-B [76], which comprises 86 ill-conditioning AutoML instances; c) UAV [77], 56 diverse conflict-free UAV path planning scenarios featured by implicit constraints multiplier in objective space (see Appendix E.3 for detail). We illustrate in Figure 3 the average optimization curves of all baselines, which is averaged within each tag and across 51 independent runs. The results show that: 1) DesignX generally shows superior optimization behavior to human-crafted optimizers from different decades, designing desirable optimizers robustly for diverse realistic problems it never saw during training; 2) DesignX consistently outperforms MetaBBO approaches, which demonstrates the novelty of our proposed bi-agent algorithm design system. By integrating two RL agents for both algorithmic workflow generation and hyper-parameter control, DesignX achieves better superior generalization performance to those MetaBBO baselines for single sub-task.
4.2 What has DesignX Learned?
Figure 4: Left: Normalized importance factors of different module types for various problem characteristics. Right: Two look-into cases for interpreting design pattern learned by DesignX.
Insightful Design Skills (RQ2). Before delving into the analysis, we first abbreviate the 10 module types in Modular-EC to simplify the presentation: Initialization (“I”), Niching (“N”), Mutation (“M”), Crossover (“C”), Other_Update (“OU”), Boundary_Control (“BC”), Selection (“S”), Population_Reduction (“PR”), Restart_Strategy (“RS”) and Information_Sharing (“IS”). The following analysis aims to investigate design principles DesignX has learned based on statistics gathered from the optimizer workflows generated for the 3200 problem instances in $\mathcal{D}_{test}$. We list several key observations we found as below:
1) In the left of Figure 4, we summarize the relative importance of different module types in Modular-EC when considering various optimization problem characteristics: Dimension, maxFEs, Search Range, Modality, Global Structure, Conditioning. To compute the relative importance, we provide a example here. Suppose we consider the relative importance of “M” (mutation) for Modality, we first divide problem instances in $\mathcal{D}_{test}$ into those unimodal ones and those multimodal ones. Then based on the optimizer workflows generated by DesignX for these problem instances, the relative importance can be calculated as the KL-divergence of the sub-module occurence distributions of “M” in unimodal problems and multimodal problems (see Appendix E.4 for more clarification). The relative importance factor reflects how DesignX thinks when designing an optimizer for a problem with certain property. As shown in Figure 4: a) for problems with different modalities, DesignX leans to design different DE mutation strategies for the generated workflow; b) for problem with different search ranges, DesignX leans to focus more on the selection of “PR” (population reduction mechanism). c) DesignX thinks designing initialization strategies has very limited impact on the final performance! These unique findings are non-trivial and deserve further analysis.
2) To investigate the above novel design principles interpreted from DesignX, we further look into the concrete sub-module occurence distributions in the first two cases. We illustrate them in the right of Figure 4. The results could clearly demonstrate DesignX’s intelligent design policy: a) for unimodal problem, it smartly choose greedy-fashion mutation operators to reinforce the optimizer’s exploitation, and dictates a composite mutation strategy for multimodal problems to address exploration and exploitation tradeoff. b) population reduction is an effective mechanism to upgrade an optimizer’s local search ability. DesignX thinks for problems with relatively smaller searching range, population reduction should be applied to accelerate the convergence. c) we examine the finding of DesignX on Initialization by replacing the designs in existing optimizers with different ones. The results validate the correctness of DesignX and is shown in Appendix F.1.
3) Another interesting design principle of DesignX is its unique taste on different optimizer types (DE, PSO, GA, ES). To illustrate this, we count the number of optimizers generated by DesignX which contain module variants derived from these four optimizer types, and then present their distribution in Figure 7. The results indicate that the DE-related algorithm sub-modules is primarily considered by DesignX to achieve aforementioned robust optimization performance. We provide several novel and very competitive DE optimizers discovered by DesignX in Appendix F.2.
Figure 5: Ratios of selected module types.
Figure 6: Averaged performance of ablation baselines.
Figure 7: Performance comparison across model sizes and training sizes.
4.3 In-depth Analysis
Ablation Study (RQ3). DesignX automates BBO optimizer design through the cooperation between Agent-1 and Agent-2. We hence investigate to what extent the two agents contribute to DesignX’s final performance. Concretely, we introduce three ablations: 1) w/o A1+A2: randomized Agent-1 & 2 without training; 2) w/o A1: only Agent-2 is trained; and 3) w/o A2: only Agent-1 is trained. We present the reversed normalized objective values (higher is better) of the ablations and DesignX on $\mathcal{D}_{test}$ and three realistic problem sets in Figure 7. Detailed results for each problem set are provided in Appendix F.3. The results reveal following insights: 1) we could at least conclude that generating a correct optimizer workflow might be more important than controlling the hyper-parameters (w/o A2 v.s. w/o A1); 2) By training DesignX via our proposed cooperative learning objective, it achieves better performance than sub-task agent, which further validates the effectiveness of our method.
Scaling Law (RQ4). We further investigate the scalability of DesignX in terms of model capacity and training data scale. Due to our limited computational resources, a preliminary study is conducted here. Specifically, we investigate three different model sizes: 1,6 and 12 layers GPT-2 blocks for both Agent-1 and Agent-2, and five training problem set sizes: 500, 2000, 5000, 10000 and 20000. We train DesignX under the corresponding 15 combinations and report their testing performance on $\mathcal{D}_{test}$ in Figure 7. y-axis denoted the average learning objective across all tested problem instances and 51 independent runs. In general, we observe that when problem set scale is small, lager model might encounter overfitting issues hence underperforms on unseen problems. In contrast, for training-instance-rich scenario, larger model’s learning ability continuously scales, while smaller ones might suffer from low capacity. However, in practice, it might consumes exponentially more resources for stable training in large models and training scales, hence in this paper, we select DesignX with 1 layer and 10k training scale as the final model. We additionally provide a comparison of DesignX and popular LLMs in terms of their design ability in Appendix F.4.
5 Conclusion
In this paper, we propose DesignX as the first end-to-end MetaBBO approach which presents human-competitive end-to-end designing ability for BBO problems. We propose a novel dual-agent system with two RL agents for optimizer workflow generation and hyper-parameters control respectively. To effectively train DesignX, we construct a large scale synthetic problem set with 10k optimization problems with diverse characteristics. A cooperative learning objective is used to harmoniously learn optimal design policies for the two RL agents. Surprisingly, a DesignX model with merely two simple GPT-2 blocks continuously surpass popular human-crafted designs along the training. We have validated the generalization ability of DesignX on both synthetic and challenging realistic scenarios. More importantly, non-trivial design principles learned by DesignX are interpreted, which provides valuable design insights back to the community. We believe DesignX could serve as a pivotal step towards fully end-to-end foundation models for automated algorithm design.
Appendix A Modular-EC
Figure 8: Left: The hierarchical Python inheritance in Modular-EC to support intricate polymorphism in EC modules. Right: Legal/Illegal algorithm workflow examples in Modular-EC.
Hierarchical Inheritance. As illustrated in the left of Figure 8, Modular-EC is designed as a Polymorphism system via multiple levels of Python inheritance. Such design allows maintaining diverse EC modules (the bottom ones in Figure 8) via universal interface encapsulation. In specific, Modular-EC holds three levels of inheritances: 1) High-level: All modules in Modular-EC stem from the abstract base class Module, which declares properties and methods shared by all modules. In high-level inheritance, two sub-classes inherit from Module: Uncontrollable and Controllable. These two sub-classes divide all possible EC modules into those without/with hyper-parameters. For Controllable, we add a config_space property as its hyper-parameter space, which for now is void until a concrete EC module is created at the low-level inheritance; 2) Middle-level: We have summarized several major EC module types from existing literature, which are widely adopted in many EC optimizers. In this inheritance level, Uncontrollable and Controllable are further divided into these EC module types. Considering that a legal (or rational) EC optimizer workflow should comprises correctly ordered modules, we add and specify a topology_rule property for each module type to indicate which module types could be placed right after it. topology_rule plays a key role in DesignX’s dual-agent algorithm design system to ensure legal generation of optimizer workflow in auto-regressive fashion. 3) Low-level: In low-level inheritance, the concrete variants of each EC module types are created, which are collected by us from existing EC literature where they serve as common choices for many EC optimizers. For a concrete low-level module variant, we assign it a unique id property as its identifier in Modular-EC, specify config_space as a dictionary of its all hyper-parameters (if it inherits from Controllable), and re-write exec() method by how it processes the solutions during optimization.
Table 2: The list of the practical variants of Controllable and Uncontrollable modules.
(a) The Controllable modules.
type Sub-module
Name + Id Functional Description Configuration Space Topology Rule
Mutation
DE/rand/1 [6]
1 - 000001 - 000000001
Generate solution $x_{i}$’s trail solution $v_{i}=x_{r1}+F1\cdot(x_{r2}-x_{r3})$
where $x_{r\cdot}$ are randomly selected solutions.
$F1\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
DE/rand/2 [6]
1 - 000001 - 000000010
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{r1}+F1\cdot(x_{r2}-x_{r3})+F2\cdot(x_{r4}-x_{r5})$
where $x_{r\cdot}$ are randomly selected solutions.
$F1,F2\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
DE/best/1 [6]
1 - 000001 - 000000011
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{best}+F1\cdot(x_{r1}-x_{r2})$
where $x_{r\cdot}$ are randomly selected solutions and $x_{best}$ is the best solution.
$F1\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
DE/best/2 [6]
1 - 000001 - 000000100
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{best}+F1\cdot(x_{r1}-x_{r2})+F2\cdot(x_{r3}-x_{r4})$
where $x_{r\cdot}$ are randomly selected solutions and $x_{best}$ is the best solution.
$F1,F2\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
DE/current-to-best/1 [6]
1 - 000001 - 000000101
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{i}+F1\cdot(x_{best}-x_{i})+F2\cdot(x_{r1}-x_{r2})$
where $x_{r\cdot}$ are randomly selected solutions and $x_{best}$ is the best solution.
$F1,F2\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
DE/current-to-rand/1 [6]
1 - 000001 - 000000110
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{i}+F1\cdot(x_{r1}-x_{i})+F2\cdot(x_{r2}-x_{r3})$
where $x_{r\cdot}$ are randomly selected solutions.
$F1,F2\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
DE/rand-to-best/1 [6]
1 - 000001 - 000000111
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{r1}+F1\cdot(x_{best}-x_{r2})$
where $x_{r\cdot}$ are randomly selected solutions and $x_{best}$ is the best solution.
$F1\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
DE/current-to-pbest/1 [65]
1 - 000001 - 000001000
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{i}+F1\cdot(x_{pbest}-x_{i})+F2\cdot(x_{r1}-x_{r2})$
where $x_{r\cdot}$ are randomly selected solutions and $x_{pbest}$ is a randomly selected from the
top $p$ best solutions.
$F1,F2\in[0,1]$, default to 0.5;
$p\in[0,1]$, default to 0.05.
Legal followers: DE-style Crossover
DE/current-to-pbest/1+archive [65]
1 - 000001 - 000001001
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{i}+F1\cdot(x_{pbest}-x_{i})+F2\cdot(x_{r1}-x_{r2})$
where $x_{r1}$ is a randomly selected solutions, $x_{r2}$ is randomly selected from the union of
the population and the archive which contains inferior solutions, $x_{pbest}$ is a randomly
selected solution from the top $p$ best solutions.
$F1,F2\in[0,1]$, default to 0.5;
$p\in[0,1]$, default to 0.05.
Legal followers: DE-style Crossover
DE/weighted-rand-to-pbest/1 [71]
1 - 000001 - 000001010
Generate solution $x_{i}$’s trail solution by $v_{i}=F1\cdot x_{r1}+F1\cdot F2\cdot(x_{pbest}-x_{r2})$
where $x_{r\cdot}$ are randomly selected solutions and $x_{best}$ is the best solution.
$F1,F2\in[0,1]$, default to 0.5;
$p\in[0,1]$, default to 0.05.
Legal followers: DE-style Crossover
DE/current-to-rand/1+archive [71]
1 - 000001 - 000001011
Generate solution $x_{i}$’s trail solution by $v_{i}=x_{i}+F1\cdot(x_{r1}-x_{i})+F2\cdot(x_{r2}-x_{r3})$
where $x_{r1},x_{r2}$ are randomly selected solutions, $x_{r3}$ is randomly selected from the union
of the population and the archive which contains inferior solutions.
$F1,F2\in[0,1]$, default to 0.5.
Legal followers: DE-style Crossover
Gaussian_mutation [5]
1 - 000001 - 000001100
Generate a mutated solution of $x_{i}$ by adding a Gaussian noise on each dimension
$v_{i}=\mathcal{N}(x_{i},\sigma\cdot(ub-lb))$ where $ub$ and $lb$ are the upper and lower bounds of the
search space.
$\sigma\in[0,1]$, default to 0.1
Legal followers: Boundary_control
Polynomial_mutation [78]
1 - 000001 - 000001101
Generate a mutated solution of $x_{i}$ as $v_{i}=\begin{cases}x_{i}+((2u)^{\frac{1}{1+\eta_{m}}}-1)(x_{i}-lb),\text{if }u\leq 0.5;\\ x_{i}+(1-(2-2u)^{\frac{1}{1+\eta_{m}}})(ub-x_{i}),\text{if }u>0.5.\end{cases}$
where $u\in[0,1]$ is a random number, $ub$ and $lb$ are the upper and lower bounds of the
search space.
$\eta_{m}\in[20,100]$, default to 20
Legal followers: Boundary_control
Multi_Mutation_1 [71]
1 - $0$ 000000001
Contains DE/current-to-pbest/1+archive, DE/current-to-rand/1+archive and DE/weighted-rand-to-best/1
three DE mutation sub-modules, its first configuration is to select one of the three mutations and the rest
configurations are to configured the selected operator.
$op\in${DE/current-to-pbest/1+archive,
DE/current-to-rand/1+archive,
DE/weighted-rand-to-best/1},
random selection in default;
$F1,F2\in[0,1]$, default to 0.5;
$p\in[0,1]$, default to 0.18.
Legal followers: DE-style Crossover
Multi_Mutation_2 [66]
1 - $0$ 000000010
Contains DE/rand/1, DE/rand/2 and DE/current-to-rand/1 three DE mutation sub-modules.
$op\in${DE/rand/1, DE/rand/2,
DE/current-to-rand/1},
random selection in default;
$F1,F2\in[0,1]$, default to 0.5;
Legal followers: DE-style Crossover
Multi_Mutation_3 [79]
1 - $0$ 000000011
Contains DE/rand/1, DE/best/2 and DE/current-to-rand/1 three DE mutation sub-modules.
$op\in${DE/rand/1, DE/best/2,
DE/current-to-rand/1},
random selection in default;
$F1,F2\in[0,1]$, default to 0.5;
Legal followers: DE-style Crossover
33 more Mutation Multi-Strategies are omitted here since they are too many
to presenting them one by one.
1 - $0$ 000000100
$\sim$1 - $0$ 000110001
$\cdots$
$\cdots$
$\cdots$
Crossover
Binomial [6]
1 - 000010 - 000000001
Randomly exchange values between parent solution $x_{i}$ and the trail solution $v_{i}$ to get a new solution:
$u_{i,j}=\begin{cases}v_{i,j},\text{ if }rand_{j}<Cr\text{ or }j=jrand\\ x_{i,j},\text{ otherwise}\end{cases},j=1,\cdots,D$ where $rand_{j}\in[0,1]$ is a
random number, $jrand\in[1,D]$ is a randomly selected index before crossover and $D$ is the
solution dimension.
$Cr\in[0,1]$, default to 0.9.
Legal followers: Boundary_control
Exponential [6]
1 - 000010 - 000000010
Exchange a random solution segment between $x_{i}$ and $v_{i}$ to get a new solution:
$u_{i,j}=\begin{cases}v_{i,j},\text{ if }rand_{k:j}<Cr\text{ and }k\leq j\leq L+k\\ x_{i,j},\text{ otherwise}\end{cases},j=1,\cdots,D$ where $k\in[1,D]$ is a randomly
selected start point for exchanging, $L\in[1,D-k]$ is a randomly determined exchange length,
$rand_{k:j}\in[0,1]^{j-k}$ is the random numbers from index $k$ to $j$ and $D$ is the solution dimension.
$Cr\in[0,1]$, default to 0.9.
Legal followers: Boundary_control
qbest_Binomial [80]
1 - 000010 - 000000011
Randomly exchange values between a solution $x^{\prime}_{i}$ selected from the top $p$ population and the trail
solution $v_{i}$ to get a new solution:
$u_{i,j}=\begin{cases}v_{i,j},\text{ if }rand_{j}<Cr\text{ or }j=jrand\\ x^{\prime}_{i,j},\text{ otherwise}\end{cases},j=1,\cdots,D$ where $rand_{j}\in[0,1]$ is a
random number, $jrand\in[1,D]$ is a randomly selected index before crossover and $D$ is the
solution dimension.
$Cr\in[0,1]$, default to 0.9;
$p\in[0,1]$, default to 0.5
Legal followers: Boundary_control
qbest_Binomial+archive [71]
1 - 000010 - 000000100
Randomly exchange values between a solution $x^{\prime}_{i}$ selected from the top $p$ population-archive union
and the trail solution $v_{i}$ to get a new solution:
$u_{i,j}=\begin{cases}v_{i,j},\text{ if }rand_{j}<Cr\text{ or }j=jrand\\ x^{\prime}_{i,j},\text{ otherwise}\end{cases},j=1,\cdots,D$ where $rand_{j}\in[0,1]$ is a
random number, $jrand\in[1,D]$ is a randomly selected index before crossover and $D$ is the
solution dimension.
$Cr\in[0,1]$, default to 0.9;
$p\in[0,1]$, default to 0.18
Legal followers: Boundary_control
SBX [81]
1 - 000010 - 000000101
Generate child solution(s) $v_{i}$ by $v_{i}=0.5\cdot[(1\mp\beta)x_{p1}+(1\pm\beta)x_{p2}]$
where $\beta=\begin{cases}(2u)^{\frac{1}{1+\eta_{c}}}-1,\text{if }u\leq 0.5;\\ (\frac{1}{2-2u})^{\frac{1}{1+\eta_{c}}},\text{if }u>0.5.\end{cases}$, $u\in[0,1]$ is a random number, $x_{p1}$ and $x_{p2}$ are two
randomly selected parents.
$\eta_{c}\in[20,100]$, default to 20
Legal followers: GA-style Mutation
Arithmetic [82]
1 - 000010 - 000000110
Generate child solution $v_{i}$ by $v_{i}=(1-\alpha)\cdot x_{p1}+\alpha\cdot x_{p2}$ where $x_{p1}$ and $x_{p2}$ are two
randomly selected parents.
$\alpha\in[0,1]$, default to 0.5.
Legal followers: GA-style Mutation
Multi_Crossover_1 [71]
1 - $0$ 000110010
Contains Binomial and qbest_Binomial+archive two DE crossover sub-modules.
$op\in${Binomial,
qbest_Binomial+archive},
random selection in default;
$Cr\in[0,1]$, default to 0.9;
Legal followers: Boundary_Control
Multi_Crossover_2 [83]
1 - $0$ 000110011
Contains Binomial and Exponential two DE crossover sub-modules.
$op\in${Binomial, Exponential},
random selection in default;
$Cr\in[0,1]$, default to 0.9;
Legal followers: Boundary_Control
9 more Crossover Multi-Strategies are omitted here since they are too many
to presenting them one by one.
1 - $0$ 000110100
$\sim$1 - $0$ 000111101
$\cdots$
$\cdots$
$\cdots$
Other_Update
Vanilla_PSO [7]
1 - 000011 - 000000001
Update solution $x^{t}_{i}$ at generation $t$ using $x^{t+1}_{i}=x^{t}_{i}+vel^{t}_{i}$ where velocity vector
$vel^{t}_{i}=w\cdot vel^{t-1}_{i}+c1\cdot rand_{1}\cdot(pbest^{t}_{i}-x^{t}_{i})+c2\cdot rand_{2}\cdot(gbest^{t}-x^{t}_{i})$,
$rand_{\cdot}\in(0,1]$ are random values, $pbest^{t}_{i}$ is the best solution $x_{i}$ ever achieved, $gbest^{t}$ is the
global best solution.
$w\in[0.4,0.9]$, default to 0.7;
$c1,c2\in[0,2]$, default to 1.49445.
Legal followers: Boundary_control
FDR_PSO [84]
1 - 000011 - 000000010
Update solution $x^{t}_{i}$ at generation $t$ using $x^{t+1}_{i}=x^{t}_{i}+vel^{t}_{i}$ where velocity vector
$vel^{t}_{i}=w\cdot vel^{t-1}_{i}+c1\cdot rand_{1}\cdot(pbest^{t}_{i}-x_{i})+c2\cdot rand_{2}\cdot(gbest^{t}-x_{i})+c3\cdot rand_{3}\cdot(nbest^{t}_{i}-x_{i})$,
$rand_{\cdot}\in(0,1]$ are random values, $pbest^{t}_{i}$ is the best solution $x_{i}$ ever achieved, $gbest^{t}$ is the global best
solution and $nbest^{t}_{i}$ is the solution that maximizes the Fitness-Distance-Ratio $nbest^{t}_{i,j}=x^{t}_{p_{j},j}$ which
$p_{j}=\mathop{\arg\max}\limits_{p\in[1,NP]}\frac{f^{t}_{i}-f^{t}_{p_{j}}}{|x^{t}_{p_{j},j}-x^{t}_{i,j}|},j=1,\cdots,D$, $f$ denotes the objective values and $D$ is solution dimension.
$w\in[0.4,0.9]$, default to 0.729;
$c1,c2\in[0,2]$, default to 1;
$c3\in[0,2]$, default to 2.
Legal followers: Boundary_control
CLPSO [64]
1 - 000011 - 000000011
Update solution $x^{t}_{i}$ at generation $t$ using $x^{t+1}_{i}=x^{t}_{i}+vel^{t}_{i}$ where velocity vector
$vel^{t}_{i}=w\cdot vel^{t-1}_{i}+c1\cdot rand_{1}\cdot(pbest^{t}_{f_{i}}-x^{t}_{i})+c2\cdot rand_{2}\cdot(gbest^{t}-x^{t}_{i})$, where $rand_{\cdot}\in(0,1]$ are
random values, $gbest^{t}$ is the global best solution, $pbest^{t}_{f_{i},j}=\begin{cases}pbest^{t}_{i,j},\text{ if }rand_{j}>Pc_{i};\\ pbest^{t}_{r,j},\text{ otherwise.}\end{cases},j=1,\cdots,D$
is the ever achieved best solution of $x_{i}$ or $x_{r}$ which is randomly selected with fitness based tournament.
$w\in[0.4,0.9]$, default to 0.7;
$c1,c2\in[0,2]$, default to 1.49445.
Legal followers: Boundary_control
CMA-ES [61]
1 - 000011 - 000000100
Given a population $x_{t}$ and the corresponding objective values $y_{t}$ at generation $t$, CMA-ES updates its
Gaussian mean $\omega_{t}$, covariance matrix $C_{t}$, and global step size $\sigma_{t}$ following [61], then samples the next
population $x_{t+1}\sim N\left(\omega_{t},\sigma_{t}^{2}\cdot C_{t}\right)$
$cc\in[0.1,1]$, default to 1;
$cs\in[0.1,1]$, default to 1.
Legal followers: Boundary_control
Sep-CMA-ES [85]
1 - 000011 - 000000101
Given a population $x_{t}$ and the corresponding objective values $y_{t}$ at generation $t$, Sep-CMA-ES updates its
Gaussian mean $\omega_{t}$, diagonal elements for the covariance matrix $D_{t}$, and global step size $\sigma_{t}$ following [85],
then samples the next population $x_{t+1}\sim N\left(\omega_{t},\sigma_{t}^{2}\cdot D_{t}\right)$
$cc\in[0.1,1]$, default to 1;
$cs\in[0.1,1]$, default to 1.
Legal followers: Boundary_control
MMES [73]
1 - 000011 - 000000110
By incorporating the Fast Mixture Sampling (FMS) [73] into a generic $(\mu,\lambda)$-ES, the next population
is sampled by $x^{i}_{t+1}\sim\omega_{t}+\sigma_{t}\cdot z^{i}_{t}$ where $\omega_{t}$ is the Gaussian mean, $\sigma_{t}$ is the mutation strength, and $z^{i}_{t}$ is a
mutation vector sampled by FMS.
$cc\in[0.1,1]$, default to 1;
$cs\in[0.1,1]$, default to 1.
Legal followers: Boundary_control
Multi_PSO_1 [86]
1 - $0$ 000001010
Contains FDR_PSO and CLPSO two PSO update sub-modules.
$op\in${FDR_PSO, CLPSO,},
random selection in default;
$w\in[0.4,0.9]$, default to 0.729;
$c1,c2\in[0,2]$, default to 1;
$c3\in[0,2]$, default to 2.
Legal followers: Boundary_Control
3 more Multi-Strategies about Other_Updates are omitted here since they are too many
to presenting them one by one.
1 - $0$ 000001011
$\sim$1 - $0$ 000001101
$\cdots$
$\cdots$
$\cdots$
Information_Sharing
Sharing
1 - 000101 - 000000001
Receive the best solution from the target sub-population and replace the worst solution in
current sub-population.
$target\in[1,N_{nich}]$, random selection in default
Legal followers: Population_Reduction, $end$
Summary of Maintained EC Modules. There are 6 Uncontrollable module types without hyper-parameters grouped in Modular-EC:
  1. 1.
    Initialization [41], which initialize a population of solutions to kick start a EC optimizer. We have maintained 5 initialization variants in the low-level inheritance (e.g., Sobol sampling [87], LHS sampling [88]).
  2. 2.
    Niching [44], which divides the population into several sub-populations. We have maintained 3 niching variants in the low-level inheritance (Random [89], Ranking [90] and Distance [91]).
  3. 3.
    Boundary_Control [42], which ensures that the values of solutions in the population are all controlled in the bound. We have maintained 5 boundary control variants in the low-level inheritance (e.g., Clip [42], Reflect [42]).
  4. 4.
    Selection [43], which selects better solutions from parents/offsprings. We have maintained 6 variants of this type in the low-level inheritance (e.g., DE-Crowding [72], GA-Roulette [5]).
  5. 5.
    Restart_Strategy [45], which re-initializes the population when it converges or stagnates. We have maintained 4 restart strategy variants in the low-level inheritance (e.g., Stagnation [92], Obj_Convergence [72]).
  6. 6.
    Population_Reduction [46], which reduces the population size to perform exploitative optimization. We have maintained 2 variants of this type in the low-level inheritance (Linear [93] and Non-Linear [94]).
(a) The Uncontrollable modules.
type Sub-module
Name + Id Functional Description Topology Rule
Initialization
Uniform [41]
0 - 000001 - 000000001
Uniformly sample solutions in the search range $x\sim\mathcal{U}(lb,ub)$
where $ub$ and $lb$ are the upper and lower bounds of the search space.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
Sobol [95]
0 - 000001 - 000000010
Sample population in Sobol’ sequences.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
LHS [88]
0 - 000001 - 000000011
Sample population in Latin hypercube sampling.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
Halton [96]
0 - 000001 - 000000100
Sample population in Halton sequence.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
Normal [97]
0 - 000001 - 000000101
Sample solutions in Normal distribution $x\sim\mathcal{N}((ub+lb)/2,\frac{1}{6}(ub-lb))$
where $ub$ and $lb$ are the upper and lower bounds of the search space.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
Niching
Rand [89]
0 - 000010 - 000000001
Randomly partition the overall population into $N_{nich}\in[2,4]$ same size
sub-populations.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
Ranking [90]
0 - 000010 - 000000010
Sort the population according to their fitness and partition them into
$N_{nich}\in[2,4]$ same size sub-populations.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
Distance [91]
0 - 000010 - 000000011
Randomly select a solution and assign its $NP//N_{nich}-1$ nearest solutions
to a new sub-population, until all solutions are assigned.
Legal followers: DE-style Mutation, PSO_Update,
GA-style Crossover
Boundary_Control
Clip [42]
0 - 000011 - 000000001
Clip the solutions out of bounds at the bound $x_{i}=\text{clip}(x_{i},lb,ub)$
Legal followers: Selection
Rand [42]
0 - 000011 - 000000010
Randomly regenerate those out of bounds $x_{i,j}=\begin{cases}x_{i,j},\text{ if }lb_{j}\leq x_{i,j}\leq ub_{j},\\ \mathcal{U}(lb_{j},ub_{j}),\text{ otherwise}\end{cases}$.
Legal followers: Selection
Periodic [42]
0 - 000011 - 000000011
Consider the search range as a closed loop
$x_{i,j}=\begin{cases}x_{i,j},\text{ if }lb_{j}\leq x_{i,j}\leq ub_{j},\\ lb_{j}+((x_{i,j}-ub_{j})\mod(ub_{j}-lb_{j})),\text{ otherwise}\end{cases}$.
Legal followers: Selection
Reflect [42]
0 - 000011 - 000000100
Reflect the values that hit the bound $x_{i,j}=\begin{cases}2ub_{j}-x_{i,j},\text{ if }ub_{j}<x_{i,j},\\ 2lb_{j}-x_{i,j},\text{ if }x_{i,j}<lb_{j},\\ x_{i,j},\text{ otherwise}\end{cases}$
Legal followers: Selection
Halving [42]
0 - 000011 - 000000101
Halve the distance between the $x_{i}$ and the crossed bound
$x_{i,j}=\begin{cases}x_{i,j}+0.5\cdot(x_{i,j}-ub_{j}),\text{ if }ub_{j}<x_{i,j},\\ x_{i,j}+0.5\cdot(x_{i,j}-lb_{j}),\text{ if }x_{i,j}<lb_{j},\\ x_{i,j},\text{ otherwise}\end{cases}$
Legal followers: Selection
Selection
DE-like [6]
0 - $0$ 000000001
Select the better one from the parent solution and its trail solution.
Legal followers: Restart_Strategy,
Population_Reduction, $end$,
Information_Sharing (If Niching is used)
Crowding [72]
0 - $0$ 000000010
The trail solution complete against its closest solution
and the better one survives.
Legal followers: Restart_Strategy,
Population_Reduction, $end$,
Information_Sharing (If Niching is used)
PSO-like [7]
0 - $0$ 000000011
Replace the old population with the new solutions
without objective value comparisons.
Legal followers: Restart_Strategy,
Population_Reduction, $end$,
Information_Sharing (If Niching is used)
Ranking [98]
0 - $0$ 000000100
Select solutions for the next generation according to the ranking based
probabilities, with the worst one ranking 1, the probability of the solution
rank $i$ is $p_{i}=\frac{1}{NP}(p^{-}+(p^{+}-p^{-})\frac{i-1}{NP-1})$ where $NP$ is the population
size, $p^{+}$ is the probability of selecting the best solution and $p^{-}$ is the
probability of selecting the worst one.
Legal followers: Restart_Strategy,
Population_Reduction, $end$,
Information_Sharing (If Niching is used)
Tournament [99]
0 - $0$ 000000101
Randomly pair solutions and select the better
one in each pair for the next generation.
Legal followers: Restart_Strategy,
Population_Reduction, $end$,
Information_Sharing (If Niching is used)
Roulette [5]
0 - $0$ 000000110
Select solutions according to the fitness based probabilities $p_{i}=\frac{f^{\prime}_{i}}{\sum_{j=1}^{NP}f^{\prime}_{j}}$
where $f^{\prime}_{j}$ is the fitness of the $j$-th solution and $NP$ is population size.
Legal followers: Restart_Strategy,
Population_Reduction, $end$,
Information_Sharing (If Niching is used)
Restart_Strategy
Stagnation [92]
0 - 000101 - 000000001
Reinitialize the population if the improvement of the best objective
value is equal to or less than a threshold $10^{-10}$ for 100 generations.
Legal followers: $end$
Obj_Convergence [72]
0 - 000101 - 000000010
Reinitialize the population if the maximal difference of the objective
values of the top 20% solutions is less than a threshold $10^{-16}$.
Legal followers: $end$
Solution_Convergence [100]
0 - 000101 - 000000011
Reinitialize the population if the maximal difference of the solutions
on all dimensions are less than a threshold $10^{-16}$ search space diameter.
Legal followers: $end$
Obj&Solution_Convergence [101]
0 - 000101 - 000000100
Reinitialize the population if the maximal difference of the objective
values is less than threshold $10^{-8}$ and the maximal distance among
solutions is less than 0.005 search space diameter.
Legal followers: $end$
Population_Reduction
Linear [93]
0 - 000110 - 000000001
Linearly reduce the population size from the initial size $NP_{max}$ to the
minimal population size $NP_{min}$. The size at generation $g+1$ is
$NP_{g+1}=round((NP_{min}-NP_{max})\cdot\frac{g}{H})+NP_{max}$
where $g$ is the generation number and $H$ is the optimization horizon.
Legal followers: Restart_Strategy, $end$
Non-Linear [94]
0 - 000110 - 000000010
Non-linearly determine the $g+1$ generation population size as
$NP_{g+1}=round((NP_{min}-NP_{max})^{1-g/H}+NP_{max})$
where $NP_{min}$ and $NP_{max}$ are the minimal and maximal population
sizes, $g$ is the generation number and $H$ is the optimization horizon.
Legal followers: Restart_Strategy, $end$
$end$
$end$
0 - 000111 - 000000001
A token indicating the completion of algorithm structure generation
which has no practical function.
For Controllable modules, we introduce four types:
  1. 1.
    Mutation [47], which introduces stochastic local search for each solution. We have maintained 49 mutation variants in the low-level inheritance (e.g., GA-gaussian [5], DE/rand/1 [6]).
  2. 2.
    Crossover [48], which encourages global optimization by exchanging two solution’s information. We have maintained 17 crossover variants in the low-level inheritance (e.g., GA-SBX [5], DE-binomial [6]).
  3. 3.
    Other_Update, which denotes other population update paradigm in PSO/ES variants. We have maintained 10 variants of this type in the low-level inheritance (e.g., ES-CMA [61], ES-Diagonal [85], PSO-normal [7]).
  4. 4.
    Information_Sharing [49], which takes the best solution in the target sub-population to replace the worst solution in current sub-population to perform information sharing between sub-populations.
Additionally, advanced evolutionary computation (EC) methods often integrate multiple candidate operators to dynamically select operators during optimization. To accommodate such scenarios, we introduce Multi_Strategy modules, which contains 2-5 candidate sub-modules of the same type (e.g., mutation operators) and expose an additional operator selection parameter in their configuration space (config_space). For categorization, Multi-Strategy modules inherit the type of their constituent sub-modules. For example, a Multi_Strategy module containing DE mutation operators is classified under the Mutation category.
Module’s ID. The unique identifier id of a module variant is a $16$-bit binary code of which: 1) the first bit is 00 or $1$ to denote if this variant is Uncontrollable or Controllable; 2) the $2$-nd to $7$-th bits denote which one of the 11 module types this variant belongs to; 3) the last $9$ bits denotes its id within that module type.
Module’s Topology Rule. A key property in a module variant is its topology_rule, which is a list of module types indicating which module types are allowed to be placed right after this module variant. A very simple example is illustrated in the right of Figure 8, where in a EC optimizer, selection modules are not allowed to be placed before crossover modules. We list some other examples here: 1) Any EC optimizer must start with Initialization; 2) Boundary_Control is not allowed placed between two subsequent reproduction modules (e.g., Mutation and Crossover); 3) Restart_Strategy is only allowed to be placed at the end of a EC optimizer.
In total, we have created 116 module variants in the low-level inheritance to cover commonly used techniques in existing EC literature. Besides, an $end$ token is included to indicate the end of the algorithm generation. We provide a complete information table about these module variants in Table 2(a) and Table 3(a), including their names, types, original papers and hyper-parameters (config_space). Such a comprehensive module space in Modular-EC could express BBO optimizers with diverse workflow structures, hence allows learning for effective (even optimal) algorithm design policies.
Appendix B Feature Design
B.1 ELA Features for Agent-1
In this paper for each problem we introduce a 13-dimensional feature vector $\mathcal{F}_{p}$ comprising two components: the 4-dimensional basic problem information and the 9-dimensional ELA features. The basic information includes the problem dimension ($D$), maxFEs ($maxFEs$), upper bound ($ub$) and lower bound ($lb$). Since the scale of these values could vary, we normalize the feature of problem dimension $\mathcal{F}_{D}=\frac{1}{5}\log_{10}D$ and the feature of maxFEs $\mathcal{F}_{FEs}=\frac{1}{10}\log_{10}maxFEs$. For the upper and lower bounds we use $\mathcal{F}_{ub}=ub/100$ and $\mathcal{F}_{lb}=lb/100$ respectively. For the 9-dimensional ELA features which are significant independence and efficient computation according to the sensitivity analysis of ELA features in [55, 56], we present them in Table 4. These features profile the optimization properties of the problem such as modality, Skewness, global structures, etc.
Table 4: The list of the ELA features for Agent-1.
Features Description
ela_meta.lin_simple.intercept The intercept of the linear regression model approximating the problem.
ela_meta.quad_simple.adj_r2 Adjusted coefficient of determination of the quadratic regression model without variable interactions.
ela_meta.lin_w_interact.adj_r2 Adjusted coefficient of determination of the linear regression model with variable interactions.
ic.m0 The initial partial information from the Information Content of Fitness Sequences (ICoFiS) approach [102].
ic.h_max The maximum information content from ICoFiS.
ic.eps_ratio The half partial information sensitivity from ICoFiS.
nbc.nn_nb.mean_ratio The ratio of arithmetic mean based on the distances among the nearest neighbors and the nearest better neighbors.
nbc.dist_ratio.coeff_var The coefficient of variation of the distance ratios.
ela_distr.number_of_peaks The estimation of the number of peaks in the distribution of the function values.
B.2 Statistical Features for Agent-2
The statistical feature $\mathcal{O}_{t}\in\mathbb{R}^{9}$is summarized below:
  1. 1.
    The first feature is the minimum objective value in the current (sub-)population indicating the achieved best performance of the current (sub-)population:
    $\mathcal{O}_{i,1}=\min\{\frac{f_{i}}{f^{0,*}-f^{*}}\}_{i\in[1,NP_{local}]}$ (8)
    It is normalized by the difference between the best objective value at initial optimization $f^{0,*}$ step and the global optimal objective value of the optimization problem $f^{*}$, so that the scales of the features from different tasks are in the same level. which hence stabilizes the training. $NP_{local}$ is the (sub-)population size.
  2. 2.
    The second one is the averaged normalized objective values in the current (sub-)population, indicating the average performance of the (sub-)population:
    $\mathcal{O}_{i,2}=\text{mean}\{\frac{f_{i}}{f^{0,*}-f^{*}}\}_{i\in[1,NP_{local}]}$ (9)
  3. 3.
    The variance of the normalized objective values in the current (sub-)population, indicating the variance and convergence of the (sub-)population:
    $\mathcal{O}_{i,3}=\text{std}\{\frac{f_{i}}{f^{0,*}-f^{*}}\}_{i\in[1,NP_{local}]}$ (10)
  4. 4.
    The next feature is the maximal distance between the solutions in (sub-)population, normalized by the diameter of the search space, measuring the convergence:
    $\mathcal{O}_{i,4}=\mathop{\max}\limits_{i,j\in[1,NP_{local}]}\frac{||x_{i}-x_{j}||_{2}}{||ub-lb||_{2}}$ (11)
    where $ub$ and $lb$ are the upper and lower bounds of the search space.
  5. 5.
    The dispersion difference [103] feature is calculated as the difference of the maximal distance between the top 10% solutions and the maximal distance between all solutions in (sub-)population:
    $\begin{split}\mathcal{O}_{i,5}=\mathop{\max}\limits_{i,j\in[1,10\%NP_{local}]}\frac{||x_{i}-x_{j}||_{2}}{||ub-lb||_{2}}\\ -\mathop{\max}\limits_{i,j\in[1,NP_{local}]}\frac{||x_{i}-x_{j}||_{2}}{||ub-lb||_{2}}\end{split}$ (12)
    It measures the funnelity of the problem landscape: a single funnel problem has a smaller dispersion difference while the multi-funnel landscape has larger value.
  6. 6.
    The fitness distance correlation (FDC) [104] describes the complexity of the problem by evaluating the relationship between fitness value and the distance of the solution from the optimum.
    $\mathcal{O}_{i,6}=\frac{\frac{1}{NP_{local}}\sum_{i=1}^{NP_{local}}{(f_{i}-\bar{f})(d_{i}^{*}-\bar{d}^{*})}}{\text{var}(\{d_{i}^{*}\}_{i\in[1,NP_{local}]})\cdot\text{var}(\{f_{i}\}_{i\in[1,NP_{local}]})}$ (13)
    where the $\bar{f}$ is the averaged objective value in (sub-)population, $d_{i}^{*}=||x_{i}-x^{*}||_{2}$ is the distance between $x_{i}$ and the best solution $x^{*}$, $\bar{d}^{*}=\text{mean}\{d_{i}^{*}\}_{i\in[1,NP_{local}]}$ is the averaged distance,
    var($\cdot$) is the variance.
  7. 7.
    The found global best objective among all (sub-)populations, indicating the achieved best performance of the overall optimization:
    $\mathcal{O}_{i,7}=\min\{\frac{f_{i}}{f^{0,*}-f^{*}}\}_{i\in[1,NP]}$ (14)
  8. 8.
    This feature is the FDC feature for the overall population:
    $\mathcal{O}_{i,8}=\frac{\frac{1}{NP}\sum_{i=1}^{NP}{(f_{i}-\bar{f})(d_{i}^{*}-\bar{d}^{*})}}{\text{var}(\{d_{i}^{*}\}_{i\in[1,NP]})\cdot\text{var}(\{f_{i}\}_{i\in[1,NP]})}$ (15)
  9. 9.
    The last feature is the remaining optimization budget, indicating the optimization progress:
    $\mathcal{O}_{i,9}=\frac{maxFEs-FEs}{maxFEs}$ (16)
    where $maxFEs$ is maximum allowed function evaluations and $FEs$ is the number of consumed function evaluations.
Appendix C Synthetic Problem Set Generation
Table 5: Overview of the basic problem functions.
ID Functions Modality Global Structure Conditioning
$f_{1}$ Sphere Unimodal Adequate Low
$f_{2}$ Schwefel F12 Unimodal Adequate Low
$f_{3}$ Ellipsoidal Unimodal Adequate Low
$f_{4}$ Ellipsoidal high condition Unimodal Adequate High
$f_{5}$ Bent cigar Unimodal Adequate High
$f_{6}$ Discus Unimodal Adequate High
$f_{7}$ Different Powers Unimodal Adequate High
$f_{8}$ Rosenbrock Unimodal Adequate Low
$f_{9}$ Ackley Multimodal Adequate High
$f_{10}$ Weierstrass Multimodal Weak High
$f_{11}$ Griewank Multimodal Weak Low
$f_{12}$ Rastrigin Multimodal Weak High
$f_{13}$ Buche-Rastrigin Unimodal Adequate High
$f_{14}$ Modified Schwefel Multimodal Weak High
$f_{15}$ Katsuura Multimodal Weak High
$f_{16}$ Composite Griewank-Rosenbrock Function F8F2 Unimodal Adequate Low
$f_{17}$ Escaffer’s F6 Multimodal Adequate High
$f_{18}$ Happycat Multimodal Weak Low
$f_{19}$ Hgbat Unimodal Adequate Low
$f_{20}$ Lunacek bi-Rastrigin Multimodal Weak High
$f_{21}$ Zakharov Unimodal Adequate Low
$f_{22}$ Levy Multimodal Weak High
$f_{23}$ Scaffer’s F7 Multimodal Weak Low
$f_{24}$ Step-Rastrigin Multimodal Weak Low
$f_{25}$ Linear Slope Unimodal Adequate Low
$f_{26}$ Attractive Sector Unimodal Adequate High
$f_{27}$ Step-Ellipsoidal Multimodal Weak Low
$f_{28}$ Sharp Ridge Unimodal Adequate High
$f_{29}$ Rastrigin’s F15 Unimodal Weak Low
$f_{30}$ Schwefel Multimodal Weak Low
$f_{31}$ Gallagher’s Gaussian 101-me Peaks Multimodal Weak Low
$f_{32}$ Gallagher’s Gaussian 21-hi Peaks Multimodal Weak Low
To construct the large scale synthetic problem set, we first collect 32 representative basic problem functions from popular benchmarks [58, 57], which are listed in Table 5. Given a solution $x\in\mathbb{R}^{D}$, a shift vector $o\in\mathbb{R}^{D}$ and a rotation matrix $M\in\mathbb{R}^{D\times D}$, the objective value of a $D$-dimensional basic problem with problem function $f_{b}$ is formulated as $F_{b}(x)=f_{b}(M^{\text{T}}(x-o))$. Then to enhance problem diversity, we borrow the idea from CEC benchmarks [57] and construct the “composition” and “hybrid” problems.
“composition” problems aggregate basic problems using weighted sum. It first randomly select $n$ basic problem functions as the sub-problems $\{f^{1},\cdots,f^{n}\}$ where $n\in[2,5]$. Then for the $i$-th sub-problem we generate a weight $w^{i}\in(0,1]$. Finally, the composition problem $F_{c}$ is calculated as the weighted sum of objective values of its sub-problems $F_{c}(x)=\sum_{i=1}^{n}w^{i}f^{i}(M^{\text{T}}(x-o))$ where $x$ is the solution, $o$ is the shift vector and $M$ is the rotation matrix.
“hybrid” problems decomposition solutions into several segments and evaluate these segments with different sub-problems. It first randomly decomposes $D$ problem dimensions into $n\in[2,5]$ segments with each segment $s^{i}=\{d^{i,0},\cdots,d^{i,D^{i}}\}$ where $d^{i,j}\in[1,D]$ is the index of the $j$-th dimension in the segment, $D^{i}$ is the length of the $i$-th segment satisfying $\sum_{i=1}^{n}D^{i}=D$. Then $n$ basic problem functions are selected as the sub-problems $\{f^{1},\cdots,f^{n}\}$ with dimensions $\{D^{1},\cdots,D^{n}\}$ respectively. The evaluation of hybrid problem $F_{h}$ is defined as $F_{h}(x)=\sum_{i=1}^{n}f^{i}((M^{\text{T}}(x-o))[s^{i}])$.
To construct the 12800 problem instances, for each problem, the problem type is randomly selected from “single” (basic problem), “composition” and “hybrid”. The problem dimension is chosen from $\{5,10,20,50\}$, the search range is sampled from $\{[-5,5],[-10,10],[-20,20],[-50,50]\}$ and the maxFEs is selected from $\{10000,20000,30000,40000,50000\}$. If the problem type is “single”, its problem function is randomly selected from the 32 basic problem functions. If the problem type is “composition” or “hybrid”, 2-5 sub-problems as well as their weights or dimension decompositions are determined. After the construction of 12800 problems, we then randomly split them into a training problem set $\mathcal{D}_{train}$ with 9600 problems and a testing problem set $\mathcal{D}_{test}$ with 3200 problems.
Appendix D Pseudo Code of Training
The cooperative training of DesignX is two-stage. Started by three initial models, the Agent-1 model $\pi_{\phi}$, Agent-2 actor $\pi_{\theta}$ and critic $v_{\psi}$, we firstly train Agent-1 and freeze Agent-2 models. For each epoch and each problem $p\in\mathcal{D}_{train}$ with dimension $D$, $100\cdot D$ solutions are sampled, evaluated and then used to calculate the ELA features $\mathcal{F}_{ELA}$ of problem $p$. Given the feature vector $\mathcal{F}_{p}$ concatenated by basic problem information and $\mathcal{F}_{ELA}$, Agent-1 auto-regressively generates the modules $\mathcal{A}_{p}$ using $\mathcal{F}_{p}$ as mentioned in Section 3.2.1 in the main paper. Controlled by the frozen Agent-2, $\mathcal{A}_{p}$ optimizes problem $p$ using $p.maxFEs$ function evaluations and obtains the accumulated reward $R_{p}$, which is then used to update $\pi_{\phi}$ in REINFORCE [59] manner. After training Agent-1, the well-trained model is frozen and its Agent-2’s turn. For each epoch and each problem $p\in\mathcal{D}_{train}$, Agent-1 generates an effective algorithm with modules $\mathcal{A}_{p}$. For each optimization step, the Agent-2 actor $\pi_{\theta}$ determines the hyper-parameters of the Controllable modules in $\mathcal{A}_{p}$ according to the current state $\mathcal{O}_{t}$. The controlled $\mathcal{A}_{p}$ optimizes $p$ for one step and obtains the next state $\mathcal{O}_{t+1}$ and reward $r_{t}$. For each $nstep$ optimization, the actor $\pi_{\theta}$ and critic $v_{\psi}$ are updated for $kepoch$ learning steps in a PPO [60] manner. The pseudo code is shown in Alg. 1. We omit the batch processing for better readability.
Input: Training problem set $\mathcal{D}_{train}$, Modular-EC $\mathcal{M}$, initial Agent-1 model $\pi_{\phi}$, Agent-2 actor $\pi_{\theta}$ and critic $v_{\psi}$.
Output: Well trained $\pi_{\phi}$, $\pi_{\theta}$ and $v_{\psi}$.
// Training for Agent-1;
Freeze $\pi_{\theta}$;
for $epoch=1$ to $Epoch$ do
       for each $p\in\mathcal{D}_{train}$ do
             Sample solutions $X_{ELA}\in\mathbb{R}^{100p.D\times p.D}$ and evaluate them $Y_{ELA}=p(X_{ELA})$;
             Obtain the ELA features $\mathcal{F}_{ELA}=\text{ELA}(X_{ELA},Y_{ELA})$;
             Get the feature vector $\mathcal{F}_{p}=$Concat$(p.D,p.maxFEs,p.ub,p.lb,\mathcal{F}_{ELA})$;
             Auto-regressively generate the optimizer $\mathcal{A}_{p}=\pi_{\phi}(\mathcal{F}_{p},\mathcal{M})$ following Section 3.2.1;
             Initial state $\mathcal{O}_{t=1}=\mathcal{A}_{p}.optimize(p),R_{p}=0$;
             while Termination condition is not met do
                   $a_{t}=\pi_{\theta}(\mathcal{O}_{t})$;
                   $\mathcal{O}_{t+1},r_{t}=\mathcal{A}_{p}.optimize(a_{t},p)$;
                   $R_{p}=R_{p}+r_{t}$;
             end while
            Update $\pi_{\phi}$ by in REINFORCE [59] manner;
       end for
end for
// Training for Agent-2;
Freeze $\pi_{\phi}$;
for $epoch=1$ to $Epoch$ do
       for each $p\in\mathcal{D}_{train}$ do
             Generate the optimizer $\mathcal{A}_{p}$ as Lines 7$\sim$10;
             Initial state $\mathcal{O}_{t=1}=\mathcal{A}_{p}.optimize(p)$;
             while Termination condition is not met do
                   for $step=1$ to $nstep$ do
                         $a_{t}=\pi_{\theta}(\mathcal{O}_{t})$;
                         $\mathcal{O}_{t+1},r_{t}=\mathcal{A}_{p}.optimize(a_{t},p)$;
                         Record transition $<s_{t},a_{t},s_{t+1},r_{t}>$;
                   end for
                  for $k=1$ to $kepoch$ do
                         Update actor $\pi_{\theta}$ and critic $v_{\psi}$ in PPO [60] manner;
                   end for
             end while
       end for
end for
Algorithm 1 The pseudo code of the training of DesignX
Appendix E Experimental Setup
E.1 Training Setup
In this paper, we set the embedding dimension $h=64$ and the number of attention head $k=4$ for both Agent-1 & 2. The number of blocks $L$ is $1$ for Agent-1 and $3$ for Agent-2. The maximum number of modules $M$ is 64 and the predefined maximum configuration size $N_{max}=12$. The training of both agents on $\mathcal{D}_{train}$ lasts for $Epoch=100$ epochs with a fixed learning rate $0.0001$. Agent-1 is trained with a batch size of $128$. During the training of Agent-2, for a batch of $64$ problems, PPO [60] method is used to update the policy and critic nets for $kepoch=3$ times for every $nstep=10$ rollout optimization steps. All experiments are run on two Intel(R) Xeon(R) 6458Q CPUs with 488G RAM. All baseline configurations align with their original papers.
E.2 Objective Value Normalization
Since the objective value scales of different problems can vary, averaging them directly is not fair, it cannot reflect the true performance of baselines. To normalize the values to the same scale, we use the best objective value found by random search $f^{*}_{p,RS}$ on problem $p$. Concretely, for problem $p$ we randomly sample $p.maxFEs$ solutions in the search range $[p.lb,p.ub]$ and take the best sampled objective value as $f^{*}_{p,RS}$. In the experiment, for the found best objective value $f^{*}_{p,b,i}$ of baseline $b$ in test run $i$ on problem $p$, we normalize it by $f_{p,b,i}^{*\prime}=\frac{f^{*}_{p,b,i}}{f^{*}_{p,RS}}$. Then we average the normalized objective values of baseline $b$ on all problem and all runs as the normalized averaged objective value in Table 1 in the main paper: $f_{b}=\frac{1}{|\mathcal{D}_{test}|\cdot 51}\sum_{p\in\mathcal{D}_{test}}\sum_{i=1}^{51}f_{p,b,i}^{*\prime}$. The similar procedure is conducted on the three realistic benchmarks. We also use a reversed normalized averaged objective value formulated as $1-f_{b}$ in the ablation study in Section 4.3.
E.3 Realistic Benchmark
  1. 1.
    Protein-Docking Benchmark [75], where the objective is to minimize the Gibbs free energy resulting from protein-protein interaction between a given complex and any other conformation. We select 28 protein complexes and randomly initialize 10 starting points for each complex, resulting in 280 problem instances. To simplify the problem structure, we only optimize 12 interaction points in a complex instance (12D problem).
  2. 2.
    HPO-B Benchmark [76] is an AutoML hyper-parameter optimization benchmark which includes a wide range of hyperparameter optimization tasks for 16 different model types (e.g., SVM, XGBoost, etc.), resulting in a total of 935 problem instances. The dimension of these problem instances range from 2 to 16. To save evaluation time, we adopt the continuous version of HPO-B, which provides surrogate evaluation functions for time-consuming machine learning tasks. We also note that HPO-B represents problems with ill-conditioned landscape such as huge flatten.
  3. 3.
    UAV Path Planning Benchmark [77] provides 56 terrain-based landscapes as realistic Unmanned Aerial Vehicle (UAV) path planning problems, each of which is 30D. The objective is to select given number of path nodes (x,y,z coordinates) from the 3D space, so the the UAV could fly as shortly as possible in a collision-free way.
E.4 Relative Importance Calculation
Taking the relative importance of mutation (“M”) modules on modality as an example, we first divide problem instances in $\mathcal{D}_{test}$ into those unimodal ones and those multimodal ones. Next we collect the mutation modules used in optimizers generated for unimodal and multimodal problems respectively. We count the occurence of each mutation sub-modules in the two mutation module collections as the histogram shown in the top right of Figure 4 in the main paper. Considering the occurence probabilities of different sub-modules in the two collections for unimodal and multimodal as two distributions, we then measure the relative importance of mutation modules to modality as the KL-divergence between the two distributions. For characteristics with more than two properties such as dimension, maxFEs and search range, we use the maximum KL-divergence among the distributions. Finally, to highlight the relative importance of different modules to the same problem characteristic, we conduct the mean-std standardization. Given the importance $\mathcal{I}_{\omega,\rho}$ of module $\omega\in\Omega$ to characteristic $\rho$, the standardized importance is $\mathcal{I}^{\prime}_{\omega,\rho}=\frac{\mathcal{I}_{\omega,\rho}-\text{mean}_{\varpi\in\Omega}(\mathcal{I}_{\varpi,\rho})}{\text{std}_{\varpi\in\Omega}(\mathcal{I}_{\varpi,\rho})}$, which is shown in the left of figure 4 in the main paper.
Appendix F Additional Experimental Results
F.1 Insightful Design Skills in Initialization
In Section 4.2 of the main paper, we observed that certain modules (e.g., Initialization) contribute minimally to optimizer performance. To validate this finding, we replace the Initialization modules in existing optimizers with five sampling methods: Uniform sampling [41], Sobol sampling [87], Latin-hypercube sampling (LHS) [88], Halton sampling [96] and Normal sampling [97]. The performance of optimizers with different Initialization modules on $\mathcal{D}_{test}$ and three realistic benchmarks are demonstrated in Figure 9. The results show that different Initialization modules have limited impact on the optimization performance, which validates the correctness of DesignX: The influence of different initialization methods might be diminished by subsequence more important optimization modules such as mutation modules.
Figure 9: The performance of optimizers with 5 different initialization modules on $\mathcal{D}_{test}$ and three realistic benchmarks.
Figure 10: Two examples of DesignX generated DE optimizers.
F.2 Examples of Generated DE Optimizers
In this section we provide two examples of the competitive DE optimizers discovered by DesignX in Figure 10. Figure 10(a) is a simple DE/current-to-rand/1/binomial optimizer with an archive for eliminated individuals. It could perform efficiency exploitative optimization on unimodal problems. Figure 10(b) is a relatively complex DE optimizer with two sub-populations split by a Distance-based Niching module which enhances the population diversity. The two sub-populations both use a mutation multi-strategy module containing 3 DE mutations: rand2, current-to-rand and current-to-best, followed by the Binomial crossover. The composite mutation modules not only address exploration and exploitation tradeoff but also provide Agent-2 more decision flexibility. Besides, linear population reduction modules are introduced to accelerate the convergence at the end of optimization. These designs make the optimizer superior in solving multimodal problems. The two examples validate the intelligence and effectiveness of DesignX.
F.3 Additional Results of Ablation baselines
In this section we demonstrate the detailed ablation study results for $\mathcal{D}_{test}$ and the three realistic benchmarks in Figure 11. The results validate that generating optimizer workflow (w/o A2) is more important than hyper-parameter control (w/o A1) in general cases. On the other hand, it is quite obvious that training Agent-1 and Agent-2 in a cooperative way results in better optimization performance. We also observe that the ablated models and the final DesignX model perform equally in HPO-B tasks, this might reveal that the generalization of DesignX on extremely ill-conditioned BBO scenarios is still limited. This might be addressed by some RL-based fine-tuning on specifically constructed ill-conditioned problem set.
Figure 11: Detailed performance of ablation baselines on $\mathcal{D}_{test}$ and three realistic benchmarks.
Table 6: Normalized averaged performance of DesignX and LLMs on synthetic and realistic problems.
GPT-4 Turbo Gemini-1.5 Deepseek-R1 DesignX
$\mathcal{D}_{test}$
2.21E-01
$\pm$7.68E-02
2.08E-01
$\pm$1.22E-01
2.31E-01
$\pm$7.26E-02
8.26E-02
$\pm$1.75E-01
Protein
Docking
9.72E-01
$\pm$2.57E-06
9.72E-01
$\pm$2.44E-06
9.71E-01
$\pm$2.51E-06
9.69E-01
$\pm$2.43E-06
HPO-B
3.78E-01
$\pm$1.89E-02
3.95E-01
$\pm$2.10E-02
4.36E-01
$\pm$1.98E-02
3.44E-01
$\pm$1.85E-02
UAV
1.28E-01
$\pm$1.20E-02
1.31E-01
$\pm$1.79E-02
1.25E-01
$\pm$1.23E-02
1.17E-01
$\pm$2.30E-02
F.4 Comparison to LLMs
We would like to note that Large Language Models (LLMs) is also capable of designing algorithms for diverse tasks [18, 19]. In the context of Optimization, however, the potential and expertise level of existing general LLMs may not be very ideal. To demonstrate this, in this section, we consider three LLM baselines: GPT-4 Turbo [105], Gemini-1.5 [106] and Deepseek-R1 [107], and compare their algorithm design ability with our DesignX model on $\mathcal{D}_{test}$ and three realistic problem sets. For each tested problem instance we prompt the LLMs with a design requirement: “You are an expert in Black-Box Optimization, given a problem instance with following mathematical form: xxx, and given its dimension as 10D, search range as [-10, 10], optimization budget as 10000 function evaluations. Please generate an optimizer with executable code for this problem. Do not generate explanations!". Then we execute their generated optimizer code to optimize the problem. The averaged results are shown in Table 6. DesignX significantly outperforms LLMs across all benchmarks. While LLMs is demonstrated with powerful general-task-solving capability, the results here clearly indicate their lacks of optimization-domain-specific knowledge. By checking the codes these LLMs generated, we found that these general LLMs are only capable of recognizing current task is an optimization task, while ignoring the specific problem characteristics behind. A direct demonstration is that they lean to generate a specific kind of optimizer: Vanilla DE, for almost all tested problem instances. In contrast, DesignX is trained specifically to tailor desired optimizers for diverse optimization problems. Through its learning from Modular-EC, valuable expert-level knowledge from human experts are effectively injected into the two agents. The cooperative large-scale training enables DesignX’s Agent-1 and Agent-2 learn optimal workflow generation policy and parameter control policy respectively, resulting in state-of-the-art performance.